### v2

Abstract
Library migration is the process of replacing a library with a similar one in a software project. Manual library migration is time-consuming and error-prone, as it requires developers to understand the Application Programming Interfaces (API) of both libraries, map equivalent APIs, and perform the necessary code transformations. Due to the difficulty of the library migration process, most of the existing automated techniques and tooling stop at the API mapping stage or support a limited set of libraries and code transformations. In this paper, we develop an end-to-end solution that can automatically migrate code between any arbitrary pair of Python libraries that provide similar functionality. Given the promising capabilities of Large Language Models (LLMs) in code generation and transformation, we use them as the primary engine for migration. Before building the tool, we first study the capabilities of LLMs for library migration on a benchmark of 314 real-world library migrations. We find that the best performing LLM, GPT-4o, can correctly perform 57% of the migrations. We also identify post-processing steps that can further improve the performance. Based on this, we develop MigrateLib, a command-line application that combines the power of LLMs, static analysis, and dynamic analysis to provide accurate library migration. We evaluate MigrateLib on 717 migration scenarios in 175 real-world Python applications that are not from our benchmark. We find that MigrateLib can migrate 32% of the migrations with complete correctness without any developer intervention. Of the remaining migrations, 73% of the migration-related changes are automatically done by the tool for more than half of the projects.
1 Introduction
Library migration is the process of replacing a library with an alternative one in a software project. The process involves identifying the library usage in the codebase, finding the corresponding Application Programming Interfaces (APIs) in the new library, and replacing the old API usage with the new API usage. This can be tedious and error-prone, especially when the library is extensively used in a large codebase Kula et al. (2018).
Researchers have proposed various techniques to automate the library migration process. Most of these techniques, however, are limited to only finding the analogous APIs (API mapping) Teyton et al. (2013); Alrubaye and Mkaouer (2018); Alrubaye et al. (2019b); Chen et al. (2019); Zhang et al. (2020); Di Rocco et al. (2025), but do not replace the old API usage in the codebase with the new one (code transformation). To find analogous APIs, some of these techniques mine existing migrations Teyton et al. (2013); Alrubaye and Mkaouer (2018); Alrubaye et al. (2019b) while others use machine learning models Chen et al. (2019); Di Rocco et al. (2025) and natural language processing techniques Ni et al. (2021). To the best of our knowledge, there are two techniques that actually perform code transformation Ni et al. (2021); Nikolov et al. (2025), but both techniques evaluate their approaches on a small set of selected library pairs without providing a tool that can be used in practice.
In recent years, Large Language Models (LLMs) have shown that they can be effective in various software engineering tasks Ozkaya (2023); Fan et al. (2023); Wang et al. (2024), including the necessary upstream tasks for library migration, such as code generation Nguyen and Nadi (2022); Peng et al. (2023); Murali et al. (2023), code comprehension Nam et al. (2024), and code transformation Wei et al. (2023); Xia et al. (2023); Zhou et al. (2023a). This suggests that LLMs may be capable of performing library migration. However, studies have also identified limitations of LLMs in other software engineering tasks, especially generating incorrect or incomplete code Fan et al. (2023). These previous efforts suggest that LLMs may be capable of performing library migration and that LLM-based migration tooling could be effective. However, we still need further empirical evidence to support this claim.
In this paper, our goal is to investigate the effectiveness of LLMs in library migration to allow informed design decisions for LLM-based migration tooling. We choose Python as the programming language for our study, as it is one of the most popular programming languages and has a rich ecosystem of libraries. Accordingly, we approach this in three phases as illustrated in Figure 1. In Phase 1, our goal is to systematically investigate the capabilities and challenges of using LLMs for library migration. We conduct an experiment using 314 migrations from PyMigBench Islam et al. (2023), a Python library migration benchmark we previously built. We experiment with three LLMs (Llama 3.1, GPT-4o mini, and GPT-4o) and measure the correctness of the migrations by comparing the LLM’s migrations with the original developer’s migrations. We find that the three LLMs can correctly migrate 26%, 49%, and 57% of the migrations, respectively. Using an alternative test-based evaluation on a subset of the projects that have unit tests, we find that 36%, 52%, and 64% of the LLM migrations maintain the same test results as the original developer migrations for the three LLMs, respectively. We also study the types of migration-related code changes LLMs excel at vs. struggle with. Overall, we find that LLMs are capable of library migration, and as expected, the more powerful LLM (GPT-4o) performs best. We also identify four major challenges that make solely relying on LLMs in a practical tool less effective: (1) The files that need migration are not always known beforehand. (2) All LLMs often return only part of the code, skipping parts of the code that do not require migration. (3) For a given project, a file that we migrate may have other files that depend on it which may require changes during migration. Specifically, if any functions are made asynchronous during the migration, all parts of the code that use those functions must also be made asynchronous. (4) The inconsistencies between the versions between the libraries used in the project and those assumed by the LLM can lead to incorrect migrations.
We find that the identified limitations from Phase 1 can be addressed by improving the prompt, some pre-processing and post-processing the LLM results. Accordingly, in Phase 2, we build a CLI tool, MigrateLib, that automates the full library migration process using LLMs and post-processing using static and dynamic program analysis. MigrateLib takes in a complete Python project with an existing test suite and automatically performs the migration for the specified library pair. By building a dynamic call graph, MigrateLib first identifies all files that require migration and sends those to the LLM via a prompt that instructs the LLM to perform the migration. MigrateLib then verifies the results of the migration by comparing the test results before and after migration. If any previously passing tests now fail, MigrateLib performs a post-processing step where it checks if the LLM skipped any code and merges it back. It also detects whether any functions were made async during the migration, and if so, finds other parts of the project that use those functions and makes them async as well. Finally, it runs the unit tests again to determine the final migration status.
PyMigBenchPhase 1 Explore LLMs for Library MigrationLlama 3.1GPT-4o miniGPT-4oPhase 2 Build MigrateLib Phase 3 Evaluate MigrateLibGPT-4oSimulated migrations314 migrations717 migrations
Figure 1: Overview of the contributions.
In Phase 3, we evaluate MigrateLib using a new purely test-based evaluation setup, as opposed to the benchmark-based evaluation we did in Phase 1. This allows us to conduct a larger-scale evaluation and also evaluate new migrations without the need for ground truth developer changes. We run MigrateLib on a set of 717 simulated migrations in 175 unique projects. We define a simulated migration as one that did not exist in the project’s history. Specifically, we find source libraries used in a project and migrate the project to an analogous target library. We ensure that the projects we choose have high code coverage (above 80%), so that the evaluation is reliable. Given that GPT-4o had the best performance in Phase 1, we conduct this evaluation only using GPT-4o. We find that MigrateLib is able to successfully complete 32% migrations without any manual intervention (i.e., all tests that pass before migration still pass after migration). To understand how much manual work is required to fix the remaining 484 migrations that the LLM left in an incomplete stage, we randomly select a subset of 43 diverse migrations and attempt to manually fix them. We measure the amount of manual work required to complete these migrations using the number of lines of code we need to manually edit. We find that we are able to fix 24 (56%) of our sampled migrations, where the manual work is only 27% of the total changes needed for the migration. The main reason we could not fix some of the remaining migrations is that the projects had complex logic that required a deep understanding of the project context, which was difficult to achieve.
Overall, our work provides a combination of empirical and engineering contributions. This paper is an extended version of our prior conference paper Islam et al. (2025) that only focused on Phase 1. Phases 2 and 3 are novel contributions of this work. To summarize, the contributions of this paper are:
  1. 1.
    An empirical study to explore the use of LLMs for library migration by comparing LLM changes to 314 benchmark developer changes. Our analysis systematically highlights the types of migration-related code changes LLMs excel at vs. struggle with.
  2. 2.
    We design and build an open-source end-to-end CLI tool, MigrateLib, that automates the full library migration process using a combination of LLMs and post-processing using static and dynamic program analysis.
  3. 3.
    We design a new evaluation setup that allows us to evaluate MigrateLib on a large scale and also on newer projects. We run an extensive test-based evaluation of MigrateLib on a set of 717 simulated migrations, and show that it works well with no or minimal manual editing for a diverse set of library pairs.
The source code of MigrateLib, the experimental data, and the detailed results of our experiments are available at https://figshare.com/account/articles/29602517.
2 Background and Terminology
In this section, we provide background information and terminology used in the rest of the paper.
2.1 Library migration
Library migration is the process of updating a software project to replace a used library with another one that provides similar functionality Teyton et al. (2012). The source library is the one being replaced, and the target library is the one that replaces it Teyton et al. (2012). One migration instance refers to a commit in a repository where a migration happened from a specific source library to a target library Islam et al. (2023), denoted using the notation source→target. For example, the commit b0607a26 in repository openstack/ironic is a retrying→tenacity migration.
A migration-related code change, or code change for brevity, is a minimal replacement of source library APIs with target library APIs that cannot be meaningfully reduced further without losing the semantics of the change Islam et al. (2023). A migration contains one or more code changes. The lines between the boxes in 2(c) show code changes. For example, segment P1 in 2(a) is replaced by segment D1 in 2(b).
We use subscripts pre, dev, and llm to denote data before migration, after developer’s migration, and after an LLM’s migration, respectively. For example, Codepre, Codedev, and Codellm denote three states of a piece of code. Changedev and Changellm denote code changes by developers and an LLM, respectively.
2.2 PyMigBench and PyMigTax
We use two of our prior artifacts: PyMigBench Islam et al. (2023) and PyMigTax Islam et al. (2024), in this work which we briefly describe next. PyMigBench is a dataset of real-world Python library migrations, which we mined from open source GitHub repositories. We use the latest version, version 2.2.5, in our experiments in this paper. The dataset has 321 migrations and 2,989 migration-related code changes. It includes a detailed description of each code change, including line numbers and API names, which facilitates our evaluation.
PyMigTax is a taxonomy of migration-related code changes Islam et al. (2024), built using the first version of PyMigBench. In our prior work, we validated its generalizability on additional third-party data. PyMigTax describes a code change based on three dimensions: (1) Program elements: the types of program elements involved in the change (e.g., function call and attribute); (2) Cardinality: how many source APIs are replaced with how many target APIs. For example, one-to-many cardinality means one source API is replaced with multiple target APIs; one-to-many, many-to-one, and many-to-many cardinalities are commonly referred to as higher-cardinality; and (3) Properties: Additional properties to describe the code change (e.g., element name change when the source and target APIs have different names.). The code change data in PyMigBench is already annotated with the PyMigTax categories. We use PyMigTax categories (shown in Table 3) to understand the types of code changes LLMs can handle. The PyMigTax paper Islam et al. (2024) has full category descriptions.
3 Phase 1: Exploring LLMs for library migration
We first systematically explore the capabilities of LLMs for library migration, while understanding how they handle different types of migration-related code changes. We first explain our setup and methods, and then present the results.
3.1 Experiment setup
We answer the following research questions in this phase:
  1. RQ-1.1
    How similar are the LLM migrations to the benchmark migrations? We consider the code changes stored in PyMigBench as the ground truth. We automatically check if the LLM was able to correctly perform all expected changes, while accounting for refactoring and alternative correct changes through manual review.
  2. RQ-1.2
    How many migrations pass unit tests? To evaluate run-time correctness of the migrated code, we use a second evaluation strategy where we run any available unit tests.
We start with the 321 migrations in PyMigBench for our experiments, then discard 2 migrations whose repositories are no longer public, and 5 migrations where either the pre-migration or post-migration file has syntax errors. We conduct our experiments with the remaining 314 migrations containing 2,910 migration-related code changes.
Listing 1: LLM Migration Prompt
The following Python code uses library <source-lib>. Migrate this code to use library <target-lib> instead. In the output, first explain the changes you made. Then, provide the modified code.
Do not make any changes to the code that are not related to migrating between these two libraries. Do not refactor. Do not reformat. Do not optimize. Do not
change coding style.
Provided code:
<premig-code>
3.1.1 Performing migration
We use three off-the-shelf LLMs – Llama 3.1-70B-Instruct Dubey et al. (2024), GPT-4o-mini-2024-07-18, and GPT-4o-2024-08-06 OpenAI (2024) – to perform migrations on the PyMigBench dataset. We refer to these models as Llama, Mini, and 4o in the rest of the paper, respectively.
For each file having code changes in a given migration, we use the LLM prompt template in 1 to migrate the code. The prompt is designed to ensure that the LLM performs focused, controlled migrations while maintaining transparency. By asking the LLM to explain its changes, we ensure that the LLM justifies its modifications, which can be useful for a human evaluator reviewing the migration. The prompt also aims to restrict the LLM from making unrelated modifications such that we remain focused on the task of library migration.
Since LLMs are non-deterministic Atil et al. (2024), we use a temperature of 0 and run each migration 10 times for each model. To understand the extent of variability in the generated code, we calculate the difference between the runs. We use git-diff to compute a diff between each pair of the 10 runs and normalize the number of different lines by dividing it by the total number of lines in Filepre. We find that on average, two runs on the same file are only 6.6%, 4.0% and 4.8% different for Llama, Mini, and 4o, respectively. Given this low variability and the high effort involved in manual validation, we randomly select one run for each model to evaluate the LLMs’ results. Our artifact contains all LLM migrations for all 10 runs.
3.1.2 Matching LLM migrations with developer migration
In RQ-1.1, we compare the LLM migrations with the developer migrations in PyMigBench. We do this for each code change, and then aggregate the results at the migration level, while also condidering the additional non-migration-related changes made by the developer and the LLM. We now explain the process using the example in 2(c) that shows a migration from the library requests to aiohttp.
(a) Before migration (Filepre)
(b) Developer’s migration (Filedev)
(c) LLM Migration (Filellm)
Figure 2: A sample requests→aiohttp migration.
The lines between the boxes show matching code changes.  denotes automatic matching, while  denotes manual matching. Dashed lines denote changes that are not recorded in PyMigBench. The blue check marks denote correct changes, while the orange crosses denote incorrect changes.
A. Auto change matching
We first attempt to automatically match as many LLM changes (Changellm) as possible with developer change (Changedev), while ensuring precision i.e., not falsely marking an incorrect Changellm as correct. Therefore, we consider only exact syntactic matches between Changedev and Changellm, ignoring formatting differences. We find exact matches by identifying the AST nodes related to Changedev in Filedev and then matching them to corresponding nodes in Filellm. If there are potentially multiple matches, we check the containing function and use proximity based heuristics demonstrated below. We ensure that the matched node sets translate to the same code string using ast.unparse.
Consider Changedev D1 in 2(c) which uses ClientSession() and get() from aiohttp. In Filellm, these APIs appear in multiple locations, but only Lines 15, 16, 21, and 22 are in the same containing function fetch_flights as Changedev. Based on line number proximity, we identify two node sets L2 and L4 as potential matches for this Changedev. By comparing the code strings, we find that L2 exactly matches Changedev. By the end of the auto change matching step, we get a set of matched pairs of Changedev and Changellm, a set of unmatched Changedev (uChangedev), and a set of unmatched Changellm (uChangellm).
B. Manually reviewing unmatched changes
The automatic matching ensures that no change is incorrectly matched, but it may miss some correct matches due to syntactic differences, or alternative correct implementations. The manual review step focuses on reviewing uChangedev and uChangellm. We use the online API documentation as well as the source code of the source and target libraries to understand the correct API usage in the context of the file. For 2(c), we manually find that L4 is a correct alternative to D2, noting that the LLM omitted the default value (True) of the allow_redirects argument, unlike the developer. L8 is a correct replacement of D3, where the LLM performed a refactoring. However, we find that the LLM used the function get instead of post at line 33, making L0 an incorrect alternative of D4.
After trying to match all changes in Changedev, we may find that some Changellm remain unmatched. Our aim is to categorize these as refactoring (not altering program semantics) or non-refactoring (potentially affecting code behavior). If a Changellm is refactoring, we remove it from uChangellm, as it does not impact migration correctness. In the example, the LLM introduced handling of a ValueError (L3), which changes the program behavior, so we keep it in uChangellm. Similarly, swapping variables in L5 is also non-refactoring.
Note that while PyMigBench records all migration-related code changes, it does not record additional changes indirectly related to the migration, i.e., lines that do not have a target library API usage. For example, adding async to function definitions (L1, L6, and L9) due to async calls introduced by migration are not recorded. We remove these changes from uChangellm and mark them as correct. Overall, L3 and L5 stay in uChangellm, while L1, L6, L7, and L9 get removed.
We perform a multi-reviewer manual review to ensure the quality of the matching process. The first author first manually reviews 190 code changes from one model. Two other authors then review about half of those code changes each independently. We then measure the agreement using Cohen’s Kappa Cohen (1960) to quantify the level of agreement between the reviewers. We do not reach substantial agreement, so we discuss and resolve disagreements and update our coding guideline. The two pairs of reviewers then independently review another 111 code changes, achieving a Cohen’s Kappa score of 0.83 (almost perfect) and 0.73 (substantial)  Landis and Koch (1977). Accordingly, we proceed with only the first author reviewing the remaining changes.
C. Determining code change status
Based on the matching of Changedev and Changellm, we determine each code change status as follows (example changes refer to 2(c)).
  • Correct change: LLM’s change is correct, either exactly like the developer (e.g., D1L2), or with an alternative API (e.g., D2L4), or with same APIs but with some refactoring (e.g., D3L8).
  • Incorrect change: The LLM incorrectly implemented the migration change: Used an incorrect API (e.g., get() instead of post() in L0), did not attempt to migrate it at all, or incorrectly removed part of a code.
D. Determining migration status
Based on the individual matched code changes, we now automatically determine an overall migration status:
  • Response failure: The LLM did not generate a Filellm for at least one Filepre (e.g., due to token limit or API timeout).
  • Syntax error: The LLM generated a Filellm for all Filepre, but at least one Filellm has syntax errors.
  • Incorrect: The LLM could not correctly migrate any of the changes marked in PyMigBench. We assign this status when all Changedev across the migration remain unmatched.
  • Partially correct: The LLM correctly migrated only some of Changedev. We assign this status when there are only some unmatched Changedev.
  • Correct with non-refactoring changes: The LLM correctly migrated all Changedev but also performed some non-refactoring changes. We assign this status when there are no unmatched Changedev in any file, but there are some remaining unmatched Changellm.
  • Correct: The LLM correctly migrated all Changedev for this migration, without any non-refactoring changes. We assign this status when there are no unmatched Changedev or Changellm left in any of the files after the matching process.
3.1.3 Evaluating LLM migrations using unit tests
In RQ-1.2, we assess the same LLM PyMigBench migrations from RQ-1.1, but using the unit tests available in the corresponding repositories. A test-based evaluation that runs the code ensures that the migration preserves the original behavior. We now explain the process below.
We make a copy of the repository after the developers’ migration (Codedev). We then make another copy of developer’s migration, but replace all Filedev with Filellm; we refer to this as Codellm. The idea is that Codellm is basically the version of Codedev where the LLM did the migration on behalf of the developer.
To set up a virtual environment, we need to identify the Python version used in the client repository. If this information is available in the setup.py file, we use that version. Otherwise, we resolve the version based on the migration commit date as follows. We first identify the release dates of all minor versions of Python (3.6, 3.7 etc). Then, we find the latest release date that is before the migration commit date. This is the latest Python version that was available at the time of the migration; we use this version to create the virtual environment. Next, we install the code dependencies using pyproject.toml, setup.py and requirements files. In cases where specific dependency versions are not specified in those files, we look up the version history of the dependency on PyPI and install the latest version available at the migration commit date, similar to how we resolve the Python version. This ensures that the dependencies are compatible with the code at the time of migration.
Once the virtual environment is ready, we run all the unit tests while measuring coverage in the repository on Codedev. If the run has errors, we read the error log and try to fix the errors, and run the tests again. Note that we only fix configuration or environment-related errors, not code errors. For example, if a dependency is missing, we install it; if the project requires a specific Python version, we set up the virtual environment with that version. We maintain a configuration file where we record any project-specific configurations or commands we used. We include this information in our artifact.
We find that 218 out of the 314 migrations have at least one test. Among these, 172 migrations had unresolvable errors. The failures are primarily due to missing dependencies, especially for older projects. The remaining 46 migrations have tests that we are able to successfully run on Codedev.
However, for the tests to be useful to validate migration, they must cover the migration-related code changes. Using the coverage report, we find that only 27 migrations have at least one test that covers the code changes, and among them, 25 migrations have at least one test that passes on Codedev. Thus, the evaluation in RQ-1.2 focuses only on these 25 migrations. We run the entire test suite on Codellm for these 25 migrations for each model using the same randomly selected LLM run used in RQ-1.1.
We compare the test results between Codedev and Codellm and assign the following statuses to the migrations:
  • Correct: All passing tests in Codedev also pass in Codellm.
  • Partially correct: Some of the tests that pass in Codedev pass in Codellm, but the others fail or raise run-time errors.
  • Incorrect: None of the passing tests in Codedev pass in Codellm; they all fail or raise run-time errors.
3.2 RQ-1.1 How similar are the LLM migrations to the benchmark migrations?
We now discuss the results of RQ-1.1. We first present the migration level results, then present the overall code change level results, and finally analyze the correctness of code changes by PyMigTax categories.
3.2.1 Migration level correctness
Table 1 shows the migration level results. We find that 4o performs the best with 94% of the migrations having at least one correct code change, closely followed by Mini with 93%. When considering fully correct migrations, 4o outperforms Mini considerably, with 57% compared to 49%. Llama has the lowest performance, with only 26% being fully correct and 51% having at least one correct code change. This is mainly due to a higher proportion of response failures where Llama ran into token limits. Llama also produces more syntax errors compared to the other two models (5.1% vs. 1.9% and 1.0%).
RQ-1.1: Migration level All three LLMs were able to perform correct migrations with 4o performing best: 94% of its migrations were partially correct and 57% were fully correct.
3.2.2 Overall code change level correctness
Table 1: Correctness of PyMigBench migrations (RQ-1.1, migration level)
Number (percentage) of migrations
Status Llama 3.1 GPT-4o mini GPT-4o
At least partially correct
Correct 83 (26%) 154 (49%) 179 (57%)
Correct w/ non-refactorings 28 (8.9%) 53 (17%) 55 (18%)
Partially correct 48 (15%) 84 (27%) 62 (20%)
Subtotal 159 (51%) 291 (93%) 296 (94%)
Fully incorrect
Incorrect 7 (2.2%) 6 (1.9%) 2 (0.6%)
Syntax error 16 (5.1%) 6 (1.9%) 3 (1.0%)
Response failure 132 (42%) 11 (3.5%) 13 (4.1%)
Subtotal 155 (49%) 23 (7.3%) 18 (5.7%)
Total migrations 314 (100%) 314 (100%) 314 (100%)
The 314 migrations we use for our evaluation contain a total of 2,910 code changes. However, when there is a response failure or syntax error in the LLM’s migration, we cannot evaluate the correctness of the corresponding code changes in that file. Accordingly, in this analysis level, we exclude 1,999 code changes from Llama, 944 from Mini, and 914 from 4o. This leaves us with 911, 1,966, and 1,996 code changes to evaluate for the three LLMs, respectively.
Table 2: Correctness of PyMigBench migration-related code changes (RQ-1.1, overall code change level)
Number (percentage) of code changes
Status Llama 3.1 GPT-4o mini GPT-4o
Correct 808 (89%) 1,744 (89%) 1,867 (94%)
Incorrect 103 (11%) 222 (11%) 129 (6.5%)
Evaluated code changes 911 (100%) 1,966 (100%) 1,996 (100%)
Excluded code changes 1,999 944 914
Total code changes 2,910 2,910 2,910
Table 2 shows the overall correctness of the LLMs’ migrations at the code change level. 4o performs best, with 94% of the code changes being correct. Interestingly, Mini and Llama both perform equally well at the code change level, with both having 89% correct code changes. This is in contrast to the migration level, where Mini performs better than Llama (49% vs. 26%). This suggests that while Mini attempted more code changes compared to Llama, the ones that Llama was able to attempt without failures or syntax errors were comparatively correct.
3.2.3 Code change correctness by category
Table 3: Correctness of code changes across PyMigTax categories (RQ-1.1, code change level by category)
Migs = Migrations, LPs = Library Pairs, CCs = Code Changes, #Eval = Number of code changes evaluated, %Cor = percentage of correct code changes.
PyMigTax category Frequency in PyMigBench Llama 3.1 GPT-4o mini GPT-4o
Migs LPs $C$s #Eval %Cor #Eval %Cor #Eval %Cor
Program elements
Function call 229 (71%) 105 (78%) 1,804 (60%) 518 88% 1,093 87% 1,092 93%
Import 318 (99%) 132 (99%) 732 (24%) 276 95% 580 94% 593 98%
Decorator 39 (12%) 13 (10%) 411 (14%) 197 87% 335 90% 353 87%
Attribute 54 (17%) 33 (25%) 183 (6.1%) 63 83% 103 78% 103 89%
Type 17 (5.3%) 13 (10%) 55 (1.8%) 12 75% 34 85% 34 88%
Exception 16 (5.0%) 13 (10%) 29 (1.0%) 17 65% 29 86% 29 93%
Function ref 4 (1.2%) 4 (3.0%) 12 (0.4%) 3 33% 12 58% 12 100%
Properties
NoProps 57 (18%) 36 (27%) 262 (8.8%) 43 84% 150 88% 150 97%
ElemNC 188 (59%) 88 (66%) 1,025 (34%) 398 84% 773 85% 771 90%
ArgAdd 93 (29%) 49 (37%) 578 (19%) 150 85% 244 80% 244 88%
ArgDel 66 (21%) 37 (28%) 349 (12%) 175 85% 234 87% 234 89%
ArgTrans 81 (25%) 38 (28%) 317 (11%) 115 77% 201 76% 220 77%
ParamAdd 13 (4.0%) 1 (0.7%) 127 (4.2%) 110 87% 127 95% 127 90%
ArgNC 17 (5.3%) 11 (8.2%) 112 (3.7%) 20 60% 58 59% 58 84%
AsyncTrans 13 (4.0%) 4 (3.0%) 50 (1.7%) 35 83% 50 90% 50 92%
OutTrans 16 (5.0%) 15 (11%) 49 (1.6%) 25 88% 49 78% 49 92%
Cardinality
One-to-One 200 (62%) 98 (73%) 1,552 (52%) 408 86% 877 84% 894 91%
One-to-Zero 50 (16%) 25 (19%) 297 (10%) 99 99% 290 99% 290 100%
Zero-to-One 35 (11%) 18 (13%) 198 (6.6%) 34 97% 55 87% 55 89%
Higher Card. 91 (28%) 45 (34%) 210 (7.0%) 94 70% 164 79% 164 84%
One-to-Many 52 (16%) 23 (17%) 142 (4.8%) 61 69% 103 78% 103 80%
Many-to-One 30 (9.3%) 24 (18%) 41 (1.4%) 16 62% 34 82% 34 91%
Many-to-Many 21 (6.5%) 11 (8.2%) 27 (0.9%) 17 82% 27 78% 27 89%
Total 321 134 2,989 911 89% 1,966 89% 1,996 94%
To understand if some types of code changes are more difficult to migrate, Table 3 shows the distribution of code change correctness across the different PyMigTax categories. The three columns in group “Frequency in PyMigBench” shows how often each PyMigTax category appears in all migrations (Migs), library pairs (LPs) and code changes (CCs) in PyMigBench, which provides perspective on the category’s impact in practice. This information is the same for each LLM since it describes the existing data in PyMigBench. For example, 1,804 (60%) of all 2,989 code changes involve function calls. The next three column groups show the performance of each model. Since each LLM has a different number of evaluated code changes, we show the number of evaluated code changes from each category (#Eval) alongside the proportion of these code changes that are correct (%Cor). For example, the Function call row under the Program elements group shows that Llama’s syntactically correct migrations included 518 code changes that involve function calls, and that 88% of these code changes were correctly migrated.
A. Program elements
Function calls are the most common program elements in terms of code changes (60%), and second most common in terms of migrations (71%) and library pairs (78%). All three LLMs correctly migrate a high proportion of these changes, with 88%, 87%, and 93% of the function calls being correctly migrated by the three LLMs. Almost all migrations and library pairs in PyMigBench require migrating import statements, a task that all three LLMs perform well, especially 4o correctly migrates nearly all of them (98%).
Mini slightly outperforms the other models in migrating decorators (90% vs. 87% and 87%); however, it does worst in attributes (78% vs. 83% and 89%). 4o performs better than other models in migrating types, exceptions and function references, followed by Mini. Llama and Mini both struggle with migrating function references (33% and 58% correct), while 4o correctly performs all such code changes. While this last category shows most discrepancy between the models, this program element is present in only 0.4% of the code changes, therefore does not affect the overall performance significantly.
B. Properties
The property no properties indicates that the source and target APIs are identical, and hence no changes are required. While 4o correctly leaves almost all of these APIs unchanged (97%), Llama and Mini incorrectly changed several of them, resulting in 84% and 88% correct, respectively.
Element name change is the most common property in PyMigBench. This indicates that the target APIs commonly bear different names from the source APIs. All LLMs perform well in making this change, with Llama, Mini, and 4o correctly migrating 84%, 85%, and 90% changes, respectively.
All three LLMs performed better or equally well in argument deletion compared to those with argument addition, which is expected as adding an argument requires the LLM to find a suitable replacement, while deleting an argument is a simpler task. The Argument transformation property indicates that an argument requires various changes, such as changing the type or value. The three LLMs correctly migrate only 76-77% of the argument transformation changes.
The property parameter addition to decorated function is found in migrations in just one library, click library, where some decorators require adding a parameter to the decorated function Pallets (2024). All three LLMs perform well in migrating these changes; interestingly, Mini performs better than 4o (95% vs. 90%). Mini also performed better than 4o in migrating decorators, which are common in the click library. These are the only two categories where 4o does not perform the best.
Argument name change applies when an argument representing the same semantics has different names in the source and target APIs. While 4o performs reasonably well in migrating these changes (84%), Llama and Mini fail frequently with them (only 60% and 59% correct).
C. Cardinality
One-to-One code changes are the most common in PyMigBench, and all three LLMs perform well in migrating them. One-to-Zero code changes are the ones where a source API needs to be removed, but no target needs to be added. Because it is a simple delete operation, all three LLMs migrate all or almost all of these changes correctly. Its counterpart, Zero-to-One code changes, are the ones where a target needs to be added, but no source needs to be removed. Interestingly, Llama performs the best in migrating these changes (97% correct), making this the only category where it performs better than the other models (87% and 89% correct). While previous research showed that higher-cardinality code changes are generally difficult compared to one-to-one code changes Alrubaye and Mkaouer (2018); Wang et al. (2016); Zhang et al. (2020); Huang et al. (2024), the PyMigBench dataset shows that higher cardinality changes are frequent with 34% of the library pairs requiring at least one higher-cardinality code change. Overall, we find that the LLMs are reasonably successful at migrating them, with 70%, 79%, and 84% correctness rates.
RQ-1.1: Code change level LLMs correctly migrate a high proportion of most code changes, with 4o achieving 94% correct code changes. The models perform reasonably well in migrating difficult higher-cardinality code changes (4o gets 84% correct), but struggle relatively more with argument transformation and argument name change.
3.3 RQ-1.2 How many migrations pass unit tests?
Table 4: Correctness of 25 PyMigBench migrations using their available unit tests (RQ-1.2)
Number (percentage) of migrations
Status Llama 3.1 GPT-4o mini GPT-4o
Correct 9 (36%) 13 (52%) 16 (64%)
Partially correct 1 (4.0%) 5 (20%) 2 (8.0%)
Incorrect 15 (60%) 7 (28%) 7 (28%)
Evaluated 25 25 25
Table 4 shows the results for the 25 migrations. 4o has the highest percentage of correct migrations, with 64%, followed by Mini with 52%, and Llama with 36%. Mini has more partially correct migrations than 4o and Llama, leading it to have an equal number of incorrect migrations as 4o (28%). Llama, on the other hand, has a much higher proportion of incorrect migrations (60%).
RQ-1.2 Out of 25 PyMigBench migrations with unit tests covering the migration-related code changes, the LLMs correctly migrated 36%-64%, with 4o being the highest.
3.4 Strengths of LLM for library migration
Overall, the results of this empirical study highlight the high overall correctness of LLM migrations. We now discuss the key strengths we observe.
Higher cardinality code changes
As shown in RQ-1.1, Llama, Mini, and 4o successfully migrated 70%, 79%, and 84% of the higher cardinality code changes, respectively, which the literature always viewed as complex migrations Xu et al. (2019); Alrubaye et al. (2020). Figure 3 shows an example of a many-to-many migration from twitter to tweepy done by Llama. In addition to correctly replacing the functions Twitter(), and OAuth() with OAuthHandler(), set_access_token(), and API() with correct arguments, Llama also splits the code into multiple statements, making the code more readable.
Before migration (using twitter)
4tw_api = Twitter(
5 auth = OAuth(oauth_token, oauth_secret,
6 consumer_key, consumer_secret))
7
8 ÂX
Llama’s migration (using tweepy)
4auth = OAuthHandler(
5 consumer_key, consumer_secret)
6auth.set_access_token(
7 oauth_token, oauth_secret)
8tw_api = API(auth)
Figure 3: Llama correctly migrating a many-to-many code change
Different API styles
We find that the LLMs are able to handle transformation between different API styles. For example, the library argparse provides API functions for parsing command line arguments, while click provides decorators. Figure 4 shows a migration example between these libraries. Despite the completely different API styles, the three LLMs correctly migrate 94% of code changes between this library pair.
Before migration (using argparse)
4def parse_commands():
5 p = ArgumentParser()
6 p.add_argument("src", type=Path)
7 p.add_argument("dst", type=Path)
8 p.add_argument("usedir",
9 action = "store_true")
10 return parser.parse_args()
11
12def main():
13 args = parse_commands()
14 convert(args.src, args.dst, args.usedir)
4o’s migration (using click)
4@argument("src",
5 type=click.Path())
6@argument("dst",
7 type=click.Path())
8@argument("usedir", type=bool)
9def main(src, dst, usedir):
10 convert(src, dst, usedir)
11
12
13
14ÂX
Figure 4: 4o correctly migrating between libraries that use different API styles.
Inferring changes outside of the source and target APIs
We find that LLMs can infer some additional changes beyond the immediate APIs of the source or target library. For example, the migration in Figure 5 adds the async_mode argument to the SocketIO() function. This is interesting because the API is from the Flask-SocketIO library, which is not the source (eventlet) or target (gevent) libraries. However, the SocketIO function works with several libraries, eventlet being the default Flask (2024). This means, if 4o did not explicitly set the async_mode argument to gevent, the code would break after the migration.
Before migration (using eventlet)
3socketio = SocketIO(app)
4bootstrap = Bootstrap(app)
4o’s migration (using gevent)
3socketio = SocketIO(app, async_mode="gevent")
4bootstrap = Bootstrap(app)
Figure 5: 4o performing a correct inferred change
3.5 Challenges of using LLMs for library migration
From the results of the empirical study in Phase 1 and the above strengths, we see that LLMs have strong potential for library migration. However, we also noted several challenges that would need to be addressed if LLMs are to be used for practical tooling for library migration.
  1. Challenge-1
    Files requiring migration may not be known upfront: In our experiment, we use PyMigBench to identify the files that require migration that we will send to the LLM. However, in a real-world scenario, the files requiring migration may not be known upfront. Therefore, a tool needs to first identify the files that require migration before sending them to the LLM for migration.
  2. Challenge-2
    LLMs do not always return complete code: A common issue we observe across all three LLMs is that they often do not return the complete code. Instead, they skips parts of the code, usually commenting that the next block of code remains unchanged. This is especially when the skipped part does not require migration, despite our prompt explicitly asking them to return the entire code. While this is not necessarily incorrect, as the skipped code may not require any changes, the resulting code is incomplete and cannot be provided directly to the developer.
  3. Challenge-3
    Dependents of migrated files may also require changes: In our experiment, we provide the LLMs only the files that directly use the source library, as marked in PyMigBench. However, other files that depend on the migrated files may also require changes. This is especially true when functions are made asynchronous during the migration, as all parts of the code that use those functions must also be made asynchronous.
  4. Challenge-4
    Library version compatibility: We do not mention any library version requirements in the prompt (1), and we notice that the LLMs usually pick the APIs from the latest or recent versions of the target library. In few cases, this caused runtime errors when running tests in LLM’s code, which we report as incorrect in RQ-1.2. This is because the application or the Python version used in the project may not be compatible with the latest version of the target library. In an actual migration scenario, the developers may want to migrate to a specific version of the target library that is compatible with their project.
4 Phase 2: Building MigrateLib
In Section 3, we show that LLMs can perform library migrations. However, we also identify several challenges when using LLMs for library migration, especially when building an end-to-end tool that developers can use. In this section, we describe Phase 2 of our work, where we build MigrateLib while addressing the identified challenges through pre- or post-processing steps as follows:
  1. 1.
    Before migration, MigrateLib identifies the files that use the source library, and thus require migration to address Challenge-1.
  2. 2.
    MigrateLib allows the user to specify the source and target library version, and include this in the prompt, to address Challenge-4.
  3. 3.
    After migration, MigrateLib automatically adds back any skipped code to address Challenge-2.
  4. 4.
    After migration, MigrateLib identifies the functions that become asynchronous due to the migration, and finds all other functions in the projects that use these functions, and makes them asynchronous as well to address Challenge-3.
MigrateLib works as a Command Line Interface (CLI) and takes the following as input: (1) the path to a Python project that needs library migration, (2) the source library name (the version is inferred), (3) the target library name and version, (4) the Python version in which the code should run, (5) paths to one or more requirements files, (6) and an LLM to use for the migration.
MigrateLib performs the migration in three main steps. First, it sets up a virtual environment for the project and identifies all files that require migration (Preparation step). In the next step, it prompts an LLM to migrate the identified files (LLM migration step). Finally, in the last step, it performs two post-processing steps to address Challenge-2 and Challenge-3 mentioned above to improve the migration results (Post-processing step).
We note that a key criteria for the input project is that it contains tests with high coverage, as MigrateLib uses the tests to identify the source library usage as well as to verify the migration. While the migration will run regardless of the test coverage, migration results without test validation may not be as reliable.
4.1 Preparation step
The preperation step handles three main preparation tasks needed before the actual migration:
Task 1: Set up environment
MigrateLib creates a virtual environment for the provided python version using the command python -m venv .venv, and then installs the required libraries using the pip install -r <requirements-file-path> command for each of the requirements file. It then installs the target library at the provided version using the pip install command. MigrateLib requires some additional packages, such as pytest, pytest-cov, which it also installs during this step. After the installation, MigrateLib derives the source library version by querying the installed package using the pip show <source-library-name> command.
Task 2: Run tests to get baseline
Once the environment is set up, MigrateLib runs the tests using the command python -m pytest on the project and saves the test report. We refer to this version of the test report as TestReportpremig, which is used as the baseline for the migration.
Task 3: Identify files that require migration
This task addresses Challenge-1 from Section 3, where we need to identify the files that require migration before sending them to the LLM. One obvious way a file can use a library is by importing it. For example, in Figure 6, the file main.py imports the geolib library, thus it requires migration when migrating from geolib to another library. To identify such files, MigrateLib reads and parses all Python files in the project and looks for import statements of the source library. Note that a library can be imported using a different name than the library name; a library may also have multiple import names. Therefore, MigrateLib first finds all import names of the source library using the Johnnydep library Jeantine-Glenn (2025), which extracts the import names from the wheel file of the library.
A file may also use a library API without directly importing it. An example is shown in Figure 6. Notice that the to_dict() function in the serialize.py takes as input a shape object from the geolib library, even though it does not import the library. It uses the Shape.name attribute and Shape.area() method, which are part of the geolib library. To identify such files, MigrateLib builds a run-time call graph of the project by running the tests with profiling. It then analyzes the call graph to find all files that call any function from the source library.
serialize.py
1def to_dict(shape):
2 return {
3 "name": shape.name,
4 "area": shape.area()
5 }
6 X
main.py
1from serialize import to_dict
2from geolib import Square
3import pprint
4
5sq = Square(5)
6pprint(to_dict(sq))
Figure 6: Example of using a library without importing
Overall, in Figure 6, MigrateLib will identify the file main.py requiring migration using static analysis, and serialize.py requiring migration using dynamic analysis.
4.2 LLM migration step
MigrateLib can be configured to use any LLM that adheres to OpenAI’s APIs OpenAI (2024), which has become a standard API for most models. For each file that requires migration, MigrateLib generates a prompt using the template in 2 and sends the prompt to the LLM. The prompt is based on our original prompt used in Section 4 (1), but enhanced with more specific instructions. Especially, we include the versions of the source and target libraries to address Challenge-4 from Section 3.
Listing 2: Prompt template used by MigrateLib
The following Python code currently uses the library <source-lib> version
<source-ver>. Migrate this code to use the library <target-lib> version
<target-ver> instead.
**Instructions:**
1. **Explain the Changes**: Begin the output with a brief explanation of the
specific changes you made to migrate from "<source-lib>" to "<target-lib>".
2. **Provide the Modified Code**: After the explanation, present the modified
code. Provide the entire code after migration even if only a part of it is
changed.
**Important Guidelines**:
- Only make changes directly related to migrating between "<source-lib>" and
"<target-lib>".
- Do not refactor, reformat, optimize, or alter the original coding style.
- The code given to you is part of a larger application. Do not change the names
of classes, functions, or variables, because it can break the application.
Original code:
<premig-code>
The LLM response contains the migrated code, which MigrateLib extracts and saves. When the LLM finishes migrating all files, MigrateLib replaces the original files in the project with the migrated files. It then runs the tests and generates a test report, which we refer to as TestReportllmmig.
4.3 Post-processing Step 1: Merge skipped code
We design this post-processing step to address Challenge-2 from Section 3. The goal is to add back any unchanged code that LLM skipped returning in its response. For each file that was migrated in the LLM migration step, MigrateLib generates a diff between the original file and the migrated file using Python’s difflib module. Then it attempts to detect which of the hunks were entirely removed from the original file. We make the following assumption to identify a hunk of skipped code:
  1. 1.
    The hunk should not use source library APIs such that it is not a migration-related change. Therefore, MigrateLib checks if any of the removed lines in the hunk uses any source library API. If yes, the hunk is not considered skipped.
  2. 2.
    The hunk must mostly include removal of lines. Our reasoning is that if the hunk has a similar number of additions and removals, it is likely that the hunk is replacing some code with other code, likely performing migration. If a hunk has more than 90% of removed lines compared to total added+removed lines, we consider it mostly removal.
  3. 3.
    The hunk should not be too small. This assumption avoids incorrectly considering small migration-related removals as skipped code. We consider a hunk as too small if it has less than 9 lines of code.
We determine the above criteria thresholds through experimentation on a set of migration examples. If a hunk is identified as skipped code by the above criteria, MigrateLib adds back the removed lines from the original file to the migrated file. After finishing the merging of skipped code, MigrateLib runs the tests again and generates a new test report, which we refer to as TestReportmerge.
4.4 Post-processing Step 2: Infer async transformation
This post-processing step addresses Challenge-3 from Section 3 by identifying functions that become asynchronous due to the migration, and making their transitive callers asynchronous as well. Below we describe the steps in detail using the example in Figure 7.
service.py
service_test.py
Before migration
1import requests
2
3def fetch_data(year):
4 url = f"https://api.io/data/{year}"
5 res = requests.get(url)
6 res.raise_for_status()
7 data = req.json()
8 return data
1from service import fetch_data
2
3def test_fetch_data():
4 year = 2023
5 data = fetch_data(year)
6 assert isinstance(data, dict)
7 assert "days" in data
8 assert len(data["days"]) == 365
After LLM migration
1from aiohttp import ClientSession
2
3async def fetch_data(year):
4 url = f"https://api.io/data/{year}"
5
6 async with ClientSession() as session:
7 async with session.get(url) as res:
8 res.raise_for_status()
9 data = HLCawait res.json()
10 return data
1from service import fetch_data
2
3def test_fetch_data():
4 year = 2023
5 data = fetch_data(year)
6 assert isinstance(data, dict)
7 assert "days" in data
8 assert len(data["days"]) == 365
9
10.
After Infer async transformation
1from aiohttp import ClientSession
2
3async def fetch_data(year):
4 url = f"https://api.io/data/{year}"
5
6 async with ClientSession() as session:
7 async with session.get(url) as res:
8 res.raise_for_status()
9 data = await res.json()
10 return data
1import pytest
2from service import fetch_data
3
4@pytest.mark.asyncio
5async def test_fetch_data():
6 year = 2023
7 data = await fetch_data(year)
8 assert isinstance(data, dict)
9 assert "days" in data
10 assert len(data["days"]) == 365
Figure 7: Example of applying LLM migration and then infer async transformation
The first row presents the original code before migration. The second row displays the code after LLM migration. The third row shows the after applying infer async transformation. The left column contains the main code, while the right column contains its associated test. Changes introduced by each migration step are highlighted in red.
Identify asynced functions
We call a function asynced if it was synchronous before migration, but is asynchronous after the LLM migration. For a given migrated file, MigrateLib creates an abstract syntax tree (AST) of the code before migration and another AST after migration. It then finds the function definitions that are not asynchronous in the AST before migration, but are asynchronous in the AST after migration. Since the location (line number) of the function definition may change due to the migration, MigrateLib matches the functions by their qualified name. In Figure 7, the function fetch_data is asynced.
Identify to-async functions
If a function foo calls a function bar, and bar calls a function fizz, then foo is said to transitively call bar and fizz. For a given asynced function, MigrateLib finds all its transitive callers using the call graph, and marks them as to-async, meaning that the async keyword must be added to their function definition. For example, in Figure 7, the function fetch_data in service.py is asynced after the LLM migration. This function is called by the function test_fetch_data in service_test.py, therefore, MigrateLib marks test_fetch_data as to-async.
Identify to-await calls
MigrateLib again uses the call graph to find all call sites of the asynced and to-async functions, and marks them as to-await, meaning that an await keyword must be added to these calls. For example, in the call to fetch_data in the function test_fetch_data (line 5, after LLM migration) is marked as to-await, because it is a call to an asynced function.
Upon finding all asynced functions, to-async functions, and to-await calls, MigrateLib applies the necessary code transformations using the LibCST library contributors (2025). Additionally, MigrateLib decorates any asynced and to-async test functions with the pytest.mark.asyncio decorator, and installs the pytest-asyncio package to the virtual environment, which is required to run asynchronous tests. After finishing the async transformations, MigrateLib runs the tests again and generates a new test report, which we refer to as TestReportasync.
5 Phase 3: Evaluating MigrateLib
In this phase, we want to evaluate MigrateLib on a large set of real-world projects. To enable such large-scale evaluation, we rely on tests to determine if the migration is successful. Since the majority of repositories in PyMigBench did not contain tests, we resort to the idea of simulated migrations where we find source libraries used in a project and migrate the project to use an analogous target library.
Our evaluation aims to answer the following research questions:
  1. RQ-2.1
    How many migrations can MigrateLib perform correctly? We measure how well MigrateLib performs on the simulated migrations without any manual intervention.
  2. RQ-2.2
    How much manual effort is required to complete the incomplete migrations? While the best case scenario is that MigrateLib can migrate the project without any issues, a more practical expectation is that MigrateLib would do most of the task correctly, but there are some parts that require manual intervention.
5.1 Data collection and selected model
We want to find repositories that use some source library, and then migrate them to use an analogous target library. To make the test-based evaluation reliable, we want to select repositories whose tests have good code coverage. We start by downloading a list of 95,472 Python GitHub repository metadata using SEART Dabic et al. (2021) service. We use the filters to select Python repositories that have at least one commit after January 1, 2024 to discard repositories that are not actively maintained. We also filter out repositories with less than 100 lines of code to ensure that the repositories have some code to migrate. We then filter out repositories using the following criteria before cloning them for further analysis:
  1. i.
    Large repositories: Extra-large repositories would take too long to clone and analyze, so we filter out 30,563 repositories larger than 10MB, leaving us with 64,909 repositories.
  2. ii.
    No requirements files: We use requirements files to build the virtual environment to be able to run the tests. We filter out 39,916 repositories that do not have a requirements.txt file in the root directory, leaving us with 24,993 repositories.
  3. iii.
    No test files: Python test file names usually contain the term test. We use GitHub code search API [^1] GitHub code search API: https://docs.github.com/en/rest/search/search#search-code to query files that contain the term test, and filter out a repository if it does not have any such files. We filter out 12,558 repositories that do not have any test files, leaving us with 12,435 repositories.
When a repository passes the above heuristics, we clone it and attempt to run the tests. In this process, we filter out the repositories with the following criteria:
  1. iv.
    Could not run tests: We attempt to set up a virtual environment for the repository using the requirements file and run the tests following Section 3.1.3 (Evaluating LLM migrations using unit tests). Due to various repository-specific requirements, e.g., Python version, system requirements, incomplete requirements file, we are often unable to run the tests. We filter out 10,104 repositories that we could not run the tests for, leaving us with 2,314 repositories.
  2. v.
    Low passing rate: We filter out the repositories that have less than 90% of the tests passing, which filters out 957 repositories, leaving us with 1,357 repositories.
  3. vi.
    Low coverage: We filter out repositories that have less than 80%line coverage by the tests, which filters out 557 repositories, leaving us with 800 repositories.
  4. vii.
    Tests require user interaction: We identify 2 repositories that require manual intervention and exclude them, leaving us with 798 repositories, which we use for evaluation. We store the commit SHA for the version of the repository we use to ensure reproducibility.
We then collect the dependencies of the 798 selected repositories by parsing their requirements files. This gives us an initial list of 1,421 unique source libraries. We exclude 12 libraries that are not found on libraries.io Katz (2020), leaving us with 1,409 libraries. To ensure that the libraries are commonly known, we filter out 602 libraries that have less than 100 dependents, and filter out 2 libraries that do not have a description, leaving us with 805 potential source libraries.
Listing 3: Prompt template for finding candidate analogous libraries
Identify alternative Python libraries that offer similar functionality as
<lib_name>, allowing for potential replacement in applications. The alternative libraries should be available on PyPI.
Produce the result as a JSON list, having only the library names. The produced library names should match the PyPI package names.
For each of the 805 source libraries, we find a list of candidate analogous libraries using gpt-4o-mini-2024-07-18 with the prompt template in 3. This gives us a total of 4,856 candidate library pairs. All of these pairs may not actually be analogous, i.e., they may not be suitable for migration, so we need to verify them. This manual task can be very time-consuming; therefore, we first automatically filter out some of the pairs. First, we filter out 4,215 pairs where the source library is used in less than 10 of the selected repositories. This help us to keep the manual effort of evaluating analogous pairs manageable by only selecting pairs that appear in multiple projects in our data set. Then, we filter out 192 pairs with an unmaintained potential target library. Specifically, we exclude the target libraries that have not been updated since January 1, 2024. This leaves us with 449 pairs to manually verify.
During manual validation, we look into the library documentation and confirm 60 library pairs that are indeed analogous and can be potentially migrated. We exclude the remaining because they fall into one of the following categories (i) Not analogous:Some libraries are not analogous, e.g., they provide different functionality or have different APIs. (ii) Test related libraries:For example, pytest, mock, etc. We exclude these libraries because our evaluation process relies on the tests. Therefore, migrating the test libraries themselves may break our automated evaluation. (iii) Non-API tools:Some dependencies are command line interfaces. Since they are not used in the code, we ignore them. For example, black, flake8, etc. (iv) Documentation tools:Some dependencies are used for documentation generation. Since they are not used in the code, we ignore them. For example, sphinx, mkdocs, etc.
Now that we have a list of 798 repositories and 60 analogous library pairs, we want to find which repositories actually use the source libraries in our list. It could be the case that some repositories list a requirement that is not actually used in the code. Accordingly, we parse the source code of each repository and look for import statements that use the source libraries. We find that 175 repositories use at least one of the source libraries. Note that each source library can have one or more analogous target libraries (e.g., requests→aiohttp and requests→httpx). We use the term unique migration instance to refer to the combination of a repository, a commit, a source library, and a target library. We use the latest commit of each repository at the time of our experiment (January 2025) for this purpose. We find a total of 717 unique migration instances, spanning 175 repositories and 48 library pairs that we use for our experiments. The 48 library pairs include 24 unique source libraries and 37 unique target libraries, and 55 unique libraries in total.
In our initial LLM empirical study (Section 3), we see that GPT-4o performs better than the other two models. Therefore, we use GPT-4o (gpt-4o-2024-11-20) to run MigrateLib in this evaluation.
5.2 Evaluation metrics
Recall from Section 4 that MigrateLib generates TestReportpremig before the migration, and three test reports after each migration step: TestReportllmmig after the LLM migration step, TestReportmerge after the merge skipped code, and TestReportasync after the infer async transformation. For a given post-migration test report, we define the correctness of a migration after that step as the proportion of tests that pass in the test report compared to the tests that passed in the TestReportpremig. For example, if 80 tests passed in TestReportpremig, and 60 of those tests passed in TestReportllmmig, then the correctness after the LLM migration step is 75% (60/80).
It is possible that some migrations may not be fully correct even after all the MigrateLib steps. We want to see how much migration work is still needed after MigrateLib migration to get a fully correct migration. Therefore, we randomly select one migration instance from each of the library pair, and attempt to manually fix them.
After fixing a migration as much as possible, we run the tests again and generate a new test report, which we refer to as TestReportmanual. We then compute the correctness of the manually fixed version. For the migrations which become fully correct, i.e., 100% correctness, we estimate the amount of manual work needed to fix them to answer RQ-2.2. We use line number of modified lines to estimate the amount of manual change required to complete a migration. For a migration, we compute a diff between the code before manual migration and the code after manual migration. From the diff, we calculate the number of changed lines of code by adding number of removed lines and number of added lines. We refer to this number as $\mathit{MigLOC}_{\mathit{manual}}$. Now, this number gives us the absolute number of lines changed, therefore we cannot directly compare this number across different migrations. For example, a migration that required 50 changed lines where the tool performed 48 of them and we only have to edit 2 lines should not be considered the same as a migration that required 4 lines of code and we still have to edit 2 lines. Therefore, we also measure the number of lines changed by MigrateLib during the automated migration from the diff between the code before automatic migration and the code after migration by MigrateLib. We refer to this number as $\mathit{MigLOC}_{\mathit{auto}}$. Accordingly, the total number of lines that needed to be changed for the migration is the sum of these two numbers. We use these to normalize the $\mathit{MigLOC}_{\mathit{manual}}$ as follows:
$\mathit{MigLOC}_{\mathit{manual}}^{\prime}=\frac{\mathit{MigLOC}_{\mathit{manual}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}\times 100$ (1)
Similarly, normalized automatic migration changes is calculated as the ratio of automatic changes to the total changes.
$\mathit{MigLOC}_{\mathit{auto}}^{\prime}=\frac{\mathit{MigLOC}_{\mathit{auto}}}{\mathit{MigLOC}_{\mathit{auto}}+\mathit{MigLOC}_{\mathit{manual}}}\times 100$ (2)
5.3 RQ-2.1: How many migrations can MigrateLib perform correctly?
Figure 8: Distribution of correctness of migrations.
The dashed lines indicate the mean, the solid lines indicate the mean.
Table 5: Comparison between migration steps
Metric LLM Only LLM+Merge LLM+Merge+Async
% fully correct migrations 27% 30% 32%
Median correctness 65% 80% 86%
# migrations applied to 717 260 129
# Migrations improved 88 (34%) 45 (35%)
Average improvement in correctness 46% 65%
Table 5 summarizes the results of RQ-2.1, while comparing the different migration steps. The table shows multiple metrics: the percentage of fully correct migrations is the proportion of migrations with 100% correctness. Median correctness is the median of correctness value across all evaluated migrations. Number of migrations applied to is the number of migrations where this step was applicable and thus got applied. Number of migrations improved shows the number of migrations where correctness improved in comparison to the previous step. Average improvement in correctness is the mean of improvement for the migrations where correctness improved after a post-processing step. Note that the post-processing steps are applied only when the migrations are not fully correct and when that specific post-processing step is applicable. For example, as shown in Table 5, merge skipped code is applied to 260 migrations, among which 88 (34%) migrations have improved correctness with an average improvement of 46% per migration. On the other hand, infer async transformation is applied to 129 migrations, among which 45 (35%) migrations have improved correctness with an average improvement of 65% per migration.
The frequency curve chart in Figure 8 shows the distribution of correctness of migrations after each migration step. All three distributions have concentrated values at 0% and 100%, indicating that most of the migrations either fail all tests or pass all tests. We see that the median correctness of migration increased from 65% in the LLM migration step to 80% after the merge skipped code, and further increased to 86% after the infer async transformation. Also note that the LLM-only has the highest value at 0% correctness, which decreases after each post-processing step. Similarly, the value at 100% correctness increases after each post-processing step. The distributions show that after post-processing, there are fewer migrations with zero or low correctness, and more migrations with high correctness. This indicates that the post-processing steps are effective in improving the correctness of migrations.
RQ-2.1 MigrateLib is able to fully migrate 32% of the migrations without any manual intervention. The median correctness of migrations improves from 65% with only the LLM migration step to 86% after the post-processing steps.
5.4 RQ-2.2: How much manual effort is required to complete the incomplete migrations?
RQ-2.1 shows that 233 migrations are completely correct after the automated migration process. The remaining 484 migrations will therefore require the developers to fix/complete them. To mimic what developers would do, we attempt to fix the migrations ourselves. However, the 484 migrations are between 43 unique library pairs in 147 repositories. To reduce the amount of manual work needed, while keeping the dataset diverse, we attempt to manually fix one incomplete migration from each library pair. Our selection results in a total of 43 migrations between 43 library pairs in 39 unique repositories, which we proceed to fix.
The yellow dotted line shows median automatic change across all 24 migrations. The numbers on top of each bar show the total changed lines of code (automatic + manual) for that migration.
Figure 9: Distribution of manual and automatic changes across manually fixed migrations.
We are able to fix 24 out of 43 migrations. We could not fix the remaining 19 due to one or more of the following reasons: (1) As outsiders to the project, some of the application logic was too complex for us to understand and fix the migration. (2) In one case, the LLM migrated to dataclasses instead of cattrs, which made the migration incorrect. Fixing this migration would require undoing the entire migration and re-migrating it to the correct library, which is beyond the scope of fixing an incomplete migration. (3) In some cases, the tests themselves need to be updated to account for the behavior of the new target library. In several cases, we could not figure the correct expectation in the test oracle (again due to the complexity of the application) or because the original tests had complex mocking of the source library APIs. (4) We did observe cases where additional dependent libraries require migration, beyond the source-target pair. We consider such migrations beyond scope of this work, thus we did not attempt to fix such migrations.
For the 24 migrations we were able to completely fix, we measure $\mathit{MigLOC}_{\mathit{manual}}^{\prime}$ and $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$ using Equation 1 and Equation 2, respectively. Figure 9 shows the proportion of automatic and manual changes made for each migration. The blue part of the stacked bar is the $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$, while the red part is $\mathit{MigLOC}_{\mathit{manual}}^{\prime}$. The yellow dotted line shows the median of $\mathit{MigLOC}_{\mathit{auto}}^{\prime}$  across all 24 migrations, which is 73%. This means that for a median migration, MigrateLib performed 73% of the changes, while the developers would only need to perform 27% of the changes. Note that 3 bars in the figure have no red part, this is because the manual fixes required to successfully complete the migration were not source code changes but some environment change, such as installing a missing dependency.
RQ-2.2 For the migrations that required manual fixes, MigrateLib performed a median of 73% of the migration changes, while a developers would only need to perform 27% of the changes to complete the migration.
To understand where MigrateLib fell short, we qualitatively analyze the manual fixes we had to perform. We categorize them into the following observations. We provide (adapted) examples from the actual migrations.
  • The LLM often incorrectly uses target APIs: Sometimes the target API call used by the LLM does not correctly replace the source API call. This includes incorrect import statements, incorrect arguments to API calls, missing API calls, incorrect exception types, incorrect use of async/await, missing non-dafault argument, and other API related mistakes. For example, in a click→typer migration, the LLM did not use a name for the Typer() call, which was required. The equivalent Click API does not require a name, so MigrateLib missed adding it.
  • The LLM skips part of the code: Despite explicitly mentioning in the prompt not to skip parts of the code (2), the LLM still skips part of the code in migrations, especially when the code pattern is repetitive or does not have source API usage. We observe the same behavior in Section 3, which inspired us to add the Merge skipped code post-processing step in MigrateLib to merge the skipped code with the migrated code (see Section 4.3). However, the heuristics used in Merge skipped code still fail to identify some skipped code so we had to manually add it back in the fix.
  • MigrateLib struggles with complex async migrations: Some sync to async migrations require complex changes in the code structure, such as changing class constructors to async methods, changing sync context managers to async context managers. The LLM often struggles to perform such complex changes correctly. Our infer async transformation in tool only performs simple transformations, such as adding async/await keywords and changing context managers. We manually fixed some of these complex async migrations.
  • The LLM refactors public program elements: In the LLM migration step, MigrateLib migrates each file in isolation, which can lead to incorrect migrations. For example, In the sqlalchemy→tortoise-orm migration, MigrateLib renames a public class EncryptedAlchemyBinaryField defined in the client project to EncryptedTortoiseBinaryField to match the target library’s name. This class is used in other parts of the code. Renaming the file prevents its callers from finding it, causing the errors in tests. Since MigrateLib does not pass the code that contains usage of the class in other files to the LLM, the LLM does not understand that it should not rename the class. Note that the prompt (2) does ask not to change any API names, however, the LLMs do not always strictly follow the instructions.
    Original code
    1gh_actor = os.getenv("GH_ACTOR", "nb")
    2gh_app_id = get_int_env_var("GH_APP_ID")
    MigrateLib migration
    1gh_actor = settings.get("GH_ACTOR", "nb")
    2gh_app_id = get_int_env_var("GH_APP_ID")
    Manual fix
    1settings.reload()
    2gh_actor = settings.get("GH_ACTOR", "nb")
    3gh_app_id = get_int_env_var("GH_APP_ID")
    Test code
    1@patch.dict(
    2 os.environ,
    3 {"GH_ACTOR": "testactor"}
    4)
    5def test_get_env_vars(self):
    6 expected_var = EnvVars(
    7 "testactor",
    8 )
    9 result = get_env_vars(True)
    10 actual = str(result)
    11 expected = str(expected_var)
    12 assert actual == expected
    Figure 10: Example of requiring to call an additional API in a python-dotenv→dynaconf migration.
  • The LLM does not have the dependent code: Figure 10 shows an example where the underlying LLM was not able to add an additional API call (settings.reload()) required for correct migration. This is required because the target library, dynaconf, loads the environment variables when the library is imported for the first time, and a reload is required if the environment variables are changed. The original code uses python-dotenv to load the environment variables, which reads the environment variables at runtime. Notice how the test, which is in another file that does not get passed to the LLM, updates the environment variables before running the tests. This updating happens after the library is imported, causing the tests to fail without the explicit settings.reload() call in the migrated code. If the LLM knew about the dependent code, it may be able to better understand the context of the migration.
  • The LLM is unaware of configuration requirements: In some cases, the migration requires changes in the project configuration rather than code changes. For example, in a migration from chardet→cchardet, adjusting the Python version from 3.11 to 3.9 resolved the test failures. Similarly, we made a migration from flask→sanic successful by installing an additional dependency, sanic-ext, with no changes to the migrated code. The LLM is unaware of such configuration requirements, and therefore, it cannot perform such changes.
  • Often test expectations need to be updated: There are cases where the tests require changes due to the target library’s different behavior. For example, often the source library raises a generic exception, which the tests expect, while the target library raises a more specific exception. Another example is coloroma uses integers to represent colors (green is 32), while rich uses strings (green is "[green]").
  • MigrateLib does not handle mock migrations: Some tests mock the source library’s API calls in the tests, and these mocks need to be migrated to the target library’s API calls. This is something we did not consider when designing MigrateLib, and therefore, it does not handle such cases.
  • MigrateLib does not support multi-library migrations: We design MigrateLib to migrate from just one source library to just one target library, which may not always be sufficient. The project miguelgrinberg/microblog-api is such an example, which uses flask and some other libraries that depend on flask. To fully migrate to fastapi (or any other target library), we also need to migrate all these dependent libraries.
6 Discussion
Overall, we find that MigrateLib automatically completes migration for a large portion of cases (32% migrations) and does most of the complex migration work (73% changes) for the remaining migrations, often leaving simple edits for the developers to perform (27% changes). Further improvements in migration process may potentially help in reducing the manual effort. However, the manual effort seems fairly minimal in most cases, specially for the developers who are familiar with the libraries and the project. We now discuss some potential improvement and future research and development directions for MigrateLib based on our findings.
The two post-processing steps currently included in MigrateLib shows that simple automated steps can significantly improve the correctness of migrations. However, we also discuss in Section 5.4 find that these implementations have limitations, therefore can be further improved. MigrateLib can also include a reverse-refactoring post-processing step to detect and revert unwanted refactorings. There are existing tools that can detect refactorings Tsantalis et al. (2022) and apply refactoring changes Ellis et al. (2022), which can be used to implement this step.
We provide the LLM with a fairly straight-forward context – the file that uses the source library APIs. This is often insufficient, and on the contrary, often too much if the file is large but only a small part of it needs migration. Identifying and sending more relevant part of the code may help the LLM to perform better migration. This however, comes with the cost of more pre-processing to identify the relevant code, as well as more post-processing to replace old code with the new code in the project.
Another potential improvement may be iteratively re-prompting the LLM with additional context, such as test error messages after running the tests on the migrated code. The challenge here is to identify meaningful error messages and generate a prompt that can help the LLM to understand the context better. Trying out different re-prompting strategies for this may be an interesting future research direction.
7 Threats to validity
7.1 Internal validity
There were several manual steps involved in our evaluations. Manual activities are inherently subjective and may contain errors. In Phase 1, we manually validated the correctness of code changes and labeled any non-migration related change as refactoring or non-refactoring. To minimize subjectivity, two authors independently perform these steps and resolved disagreements through discussion. In Phase 3, only one author performs all manual migrations. However, we rely on objective test outcomes as the primary measure of correctness.
The merge skipped code uses some thresholds to identify skipped code hunks, which we determine through experimentation. Different thresholds may yield different results. However, we believe our chosen thresholds strike a reasonable balance between adding back genuinely skipped code and avoiding incorrect additions. If the assumptions do not hold, the resulting migration may potentially fail, therefore, the results we have may be a lower bound of the actual migration correctness.
7.2 External validity
We use three off the shelf LLMs in Phase 1, and the best of them in Phase 3. Using an LLM specifically tuned for the library migration task could potentially yield better results. Therefore, our findings in this paper may be viewed as a lower bound of the performance of potentially more powerful and specialized models.
In Phase 3, we focus on Python projects that are well-tested and actively maintained, as selected through specific filters. This focus may limit the generalizability of our findings to projects with less test coverage, different maintenance status, or other programming languages. The set of library pairs and migration scenarios is also limited by the availability of analogous libraries and the manual validation process. As a result, the observed effectiveness of MigrateLib may not fully generalize to all migration contexts. That said, the size of the dataset (717 migrations) and the diversity of selected projects (175 repositories) provide a reasonable basis for evaluating the tool’s effectiveness in a variety of migration scenarios.
The libraries we consider in Phase 1 and Phase 3 are relatively popular and well-established. This familiarity with APIs and API usage is precisely our hope when using LLMs for migration. However, being familiar with the APIs alone does not always guarantee being able to map them to each other or apply the right transformations on all code bases, which is why our evaluation is necessary. For newer libraries, additional information, e.g., API documentation, may be needed to guide the LLM in the migration process. This can be an interesting avenue for future study, and thus improvement of MigrateLib.
7.3 Construct validity
In RQ-1.1, we analyze the code change categories based on Changedev, not on Changellm. The categories can differ if the LLM produces a different code change than the developer. PyMigBench already provides us with the Changedev categories; categorizing the LLM changes according to PyMigTax Islam et al. (2024) requires additional manual effort that is beyond the scope of this work. That said, our observation is that the code/API differences in most cases are minor, and would not yield in a different PyMigTax category.
We use automated test suites to assess migration correctness in Phase 3, but these tests may not cover all possible behaviors or edge cases in the migrated projects. To minimize this, we select projects with at least 80%lines of code covered by tests, which helps ensure that the tests are comprehensive enough to provide a reliable measure of correctness.
We use migration lines of code (Equation 1) to estimate the amount of migration work, which may not capture all aspects of migration work, such as time spent understanding or debugging the migration. We choose to use this metric since it is a measurable objective proxy for the effort required.
8 Related Work
There has been a significant amount of research aimed at automating library migration. We discuss these techniques in this section.
The majority of the existing library migration techniques focus on identifying API mappings between two analogous libraries. Some of these techniques mine existing library migrations to identify the API mappings Teyton et al. (2012, 2013); Alrubaye et al. (2019b, a, 2020). Some techniques see the requirement of existing migrations as a limitation, and therefore use other techniques, such as machine learning Alrubaye and Mkaouer (2018); Chen et al. (2019) and natural language processing Ni et al. (2021). These techniques only identify the API mappings, but do not perform client code transformation, which is a separate challenge.
SOAR Ni et al. (2021) uses API documentation to identify API mappings, and then uses the mappings and unit tests to guide program synthesis for client code transformation. SOAR has several limitations that prevent it from being a general solution for library migration. SOAR is evaluated on only two pairs of deep learning libraries, focusing on their migration of neural network models. These APIs are usually called in a sequence, where the output of one call is used as input to the next call. The nature of these APIs allows SOAR to check the program state after each API call, which is not the case for most libraries. Additionally, given the limited number of libraries, their technique relies on the library-specific error message format to further guide the synthesis process, requiring extensive changes and technique adaptations for running SOAR on other libraries. SOAR supports one-to-many code changes for a small set of pre-determined APIs, whose information is hard-coded for the transformation process. SOAR also times out for 20% of the evaluated migrations, even for small code snippets, the largest being 183 lines of code.
Zhou et al. Zhou et al. (2023b) proposed HaPiM, which first trains an unsupervised machine learning model capable of API mapping, and then uses the API mappings to guide an LLM in code transformation. Similar to Chen et al. Chen et al. (2019), HaPiM learns from API usage patterns of source and target libraries in their respective client projects, therefore does not require any existing migration examples. For the code transformation, they use the API mappings to prompt an LLM to replace source API calls with target API calls. Their approach outperforms MigrationMapper Alrubaye et al. (2019b) and GPT-3.5 Turbo on the BLEU Papineni et al. (2002) and CodeBLEU Ren et al. (2020) metrics evaluated on 5 Java library pairs. A significant limitation of HaPiM is that it works with small independent code snippets that use the source library APIs, and results in an equivalent snippet that uses the target library APIs. This is as opposed to real-life code where the source API usages are interleaved with other types of code, such as variable declarations, control statements, and other library usages. This means that to use HaPiM, developers need to first extract code snippets that contains only the source library API usages from the client code, then use HaPiM to replace them with target library APIs, and then re-integrate the transformed code back into the client code, which is not practical for large-scale library migrations.
Nikolov et al. Nikolov et al. (2025) describe Google’s experiences in using LLMs for several internal code transformation tasks, including Joda Time→Java Time migrations. They do not attempt to provide a general solution for library migration; rather, they use LLMs to perform the migration they need. Therefore, their prompts contains instructions and contextual information tailored specifically for Joda Time→Java Time migrations. Nonetheless, they show that using LLMs reduced migration time by at least 50% compared to manual migration.
There are also a few studies that focus on migrating between different versions of the same library. The study of Nikolov et al. Nikolov et al. (2025) we describe above also includes JUnit 3→JUnit 4 migrations. Almeida et al. Almeida et al. (2024) evaluate the performance of ChatGPT for SQAlchemy 1→SQLAlchemy 2 migrations. While similar, migrating between different versions of the same library is different from migrating between different libraries, because the former can leverage the evolution or release history of the libraries to identify API changes.
The above shows that there have been different approaches to automate the library migration process. However, these techniques come with one or more limitations that prevent them from being a general solution for library migration. Our tool, MigrateLib, resolves these limitations and can automate the entire library migration process for any arbitrary analogous library pair. We evaluate MigrateLib on a large and diverse dataset of 717 migrations between 48 analogous library pairs, and show that MigrateLib can successfully perform most of the migration with no or minimal manual intervention from the developer.
9 Conclusion
Our goal in this paper is to fully automate the Python library migration process. We systematically approached this goal by first conducting an empirical study to understand the capabilities of LLMs for automating Python library migration. Overall, we find that LLMs can handle complex code changes but we also identify some challenges that need to be addressed to make LLM-based migration practical. Based on these findings and insights, we built MigrateLib, a practical CLI tool that combines LLMs with static and dynamic program analysis to automate the full migration process. Our evaluation on 717 simulated migrations across 175 diverse projects and 48 library pairs shows that MigrateLib can fully automate a large portion of migrations (32%), with most remaining cases requiring only minor manual intervention (median 27% changes). The tool’s layered approach, including post-processing steps, improves LLM migration correctness (from 65% to 86%) and reduces developer effort. Overall, MigrateLib lowers the barrier for large-scale library migration in real-world projects, demonstrating accessible and effective automated library migration.
Acknowledgements.
We thank Dr. Ildar Akhmetov, Northeastern University, for his contributions to the early ideas of this work.
Declarations
Funding
This work was supported by the Canada Research Chairs Program and NSERC CREATE.
Author contributions
Below are the contributions of each author –
  • Mohayeminul Islam contributed to idea conceptualization, design and implementation of the MigrateLib, conducting the experiments, writing script for the experiments, manual evaluations in Phase 1, manual migration fix in Phase 3, and writing the manuscript.
  • Ajay Kumar Jha contributed to idea conceptualization, design of the MigrateLib, manual evaluations in Phase 1, and writing the manuscript.
  • May Mahmoud manual evaluations in Phase 1, and revising the manuscript.
  • Sarah Nadi contributed to idea conceptualization, design of the MigrateLib, manual evaluations in Phase 1, and writing the manuscript. She is the Principal Investigator of the project.
Data availability
The source code of MigrateLib, the experimental data, and the detailed results of our experiments are available at the following repository: https://figshare.com/account/articles/29602517.
Conflict of interest
No conflict.
Ethical approval/informed concent/clinical trial number
Not applicable.